# The Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future Detections

### v1

@@ 4.1 Observational data
In its numerical implementation, our model allows for simultaneous, multi-wavelength modelling. We use $94$ data points, from radio to X-rays between $9.2$ and $380$ days [^3] Data points are provided by K. Mooley on github.com/kmooley/GW170817/, as compiled and reprocessed homogeneously by Makhathini et al. (2021). We also include $5$ constraining early-time upper limits also reported in Makhathini et al. (2021), as shown in Table 4. Early-time constraints are useful to restrict the posterior models to those with a single peak in flux density, as discussed in § 2.2. All these flux measurements and upper limits are fitted simultaneously.
The likelihood that we use for our study is based on a modified $\chi^{2}$ calculation to account for the upper limits:
@@
where the first sum concerns observational data (subscripts $i$) while the second sum concerns upper limits (subscripts $j$). The quantities $m_{i}$ and $m_{j}$ are the model flux densities at each observing time and frequency of the observational data set; $d_{i}$ is the flux density value of a detection; $\sigma_{i}$ its associated uncertainty; $u_{j}$ is a flux upper limit presented in Table 4; and $\sigma_{j}$ is an equivalent uncertainty that we arbitrarily define as $\sigma_{j}=0.2u_{j}$, which roughly corresponds to the typical uncertainties for the detections. Adding this element accommodates for some models whose predictions are slightly above the reported upper limit. Note that the quantity $\max\{u_{j};m_{j}\}$ in the second term implies that it behaves as a penalty added only for predicted fluxes $m_{j}$ above the observed upper limits. Another option would be to force the likelihood to diverge as soon as one of the modelled fluxes is above the corresponding upper limit.
In this study, we choose not to include late-time observations after $400\leavevmode\nobreak\ \mathrm{days}$ in the fits. Indeed, we expect jet lateral spreading to play a role at late times, when the shock front reaches Lorentz factors $\Gamma\lesssim 3$, by which point our model cannot accurately describe the jet dynamics. In addition, other emission sites (e.g. the kilonova afterglow, Nakar & Piran 2011) could also contribute to the observed flux at late times. Including the latest observations (Hajela et al., 2022) which were not included in the compilation by Makhathini et al. (2021) is also difficult: late-time data in X-rays consists of only a few photons and a correct modelling in this case should be expressed in photon counts and account for the instrumental response and low-statistics effects. The result of such an analysis is currently unclear, with an uncertain late rebrightening (Hajela et al., 2022): see discussion in Troja et al. (2022). Interpreting the late evolution of GW 170817’s afterglow seems to require a dedicated study, and we therefore exclude this phase from the analysis presented here. Finally, we do not include the H.E.S.S. upper limit at the peak (Abdalla et al., 2020), or the Fermi/LAT and H.E.S.S. early upper limits (Ajello et al., 2018; Abdalla et al., 2017), as it does constrain the fit (see § 4.4).
@@
 [^4] Times are given in days, frequencies in Hz, and upper limits in $\mu\mathrm{Jy}$. These flux upper limits are selected as the most constraining among all the early observations listed in the compilation by Makhathini et al. (2021).
Figure 5: Comparison of the marginalized posterior distributions for the three fits of the afterglow of GW 170817 presented in Section 4, which differ only by the treatment of the radiative processes: ”no SSC” in blue; ”SSC (Thomson)” in green; ”SSC (with KN)” in orange. The first two rows show the inferred distributions of the model’s free parameters. The last row shows the distributions of several quantities derived from these parameters: true energy of the jet $E_{0}$ (see Equation (9)), ratio $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}$, and four quantities used in the conditions for single-peaked light curves from Beniamini et al. (2020b) and described in § 2.2, $\log{\left(\theta_{\mathrm{c}}\Gamma_{0}^{\mathrm{c}}\right)}$, $8b/a$, $-b\frac{\log\theta_{\mathrm{c}}}{\log\Gamma_{0}^{\mathrm{c}}}$ and $(b-1)\frac{\log{\left(\theta_{\mathrm{v}}/\theta_{\mathrm{c}}\right)}}{\log{\left(\Gamma_{0}^{\mathrm{c}}\theta_{\mathrm{c}}\right)}}$.
@@ 4.2 Results from the afterglow fitting
We perform a Bayesian analysis on the afterglow data of GW 170817 using the Markov Chain Monte Carlo (MCMC) algorithm of the Python suite emcee (Foreman-Mackey et al., 2013, 2019). We run three separate fits where only the assumptions for the radiative processes in the shocked region differ:
  1. 1.
    ”no SSC”: emission is only produced by synchrotron radiation, as described in Sari et al. (1998) and § 2.4.4;
  2. 2.
    ”SSC (Thomson)”: SSC is taken into account assuming that all scatterings occur in Thomson regime, as described in Sari & Esin (2001) and § 2.4.5;
  3. 3.
    ”SSC (with KN)”: SSC is taken into account including the KN regime, as described in Nakar et al. (2009) and § 2.4.6.
The priors used for these three fits are identical and presented in Table 3. We use (log-)uniform priors over large intervals in order to remain as agnostic as possible. Their bounds are shown in Table 3. Additional constraints on the viewing angle can be obtained by combining the GW signal and the accurate distance and localization of the host galaxy (Finstad et al., 2018), or by combining the afterglow photometry and VLBI imagery (Govreen-Segal & Nakar, 2023), and the density of the external medium can be constrained from direct observation of the host galaxy (Hallinan et al., 2017; Hajela et al., 2019). We do not include them to focus on constraints obtained from the afterglow modelling only.
Table 3: Free parameters of the MCMC, their prior bounds and shape.
@@
For each of the three fits, we initialize $50$ chains and run 10 0001000010\,00010 000 iterations per chain. After studying the convergence speed, we remove the first 2 00020002\,0002 000 iterations for each chain and show the posterior distributions for the remaining 400 000400000400\,000400 000 samples of parameters.
Table 4: Posterior values for the free parameters of the three firs of the afterglow of GW 170817.
@@
555 The three fits are presented in § 4.2. The table reports the median value of each parameter’s posterior distribution, with uncertainties corresponding to the 90% credible intervals.
We present the marginalized posterior distribution of the free parameters for the three fits in Fig. 5, complemented by Fig. 16 in appendix, which shows the corresponding joint and marginalized posterior distributions at $3\sigma$ credibility (”corner plot”). We also show in Fig. 5 the distribution of some derived quantities, the true initial kinetic energy of the ejecta $E_{0}$ as deduced from Equation (9), the ratio of the viewing angle over the opening angle of the core jet $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}$ and the four conditions for single-peaked light curves described by Beniamini et al. (2020b) and in § 2.2. The three fits lead to similar distributions across the parameter space, with a few exceptions discussed in § 4.3. The median parameter values as well as their 90% credible intervals are reported in Table 5. We finally show in Table 6 the median values of the derived quantities shown in Fig. 5.
@@
 [^6] The three fits are presented in § 4.2 and the posterior values of the free parameters are reported in Table 5. The table reports the median value of each quantity’s posterior distribution, with uncertainties corresponding to the 90% credible intervals. The four last rows correspond to the quantities appearing in the conditions for single peak light curves (§ 2.2).
The detailed joint posterior distributions for the ”SSC (with KN)” model where emission is produced by synchrotron radiation and SSC diffusions in both Thomson and KN regimes are shown in Fig. 6. In this figure, which is representative of the results obtained with the three fits, we observe typical correlations between some of the afterglow parameters, like $\theta_{\mathrm{c}}$ and $\theta_{\mathrm{v}}$, $E_{\mathrm{0,iso}}^{\mathrm{c}}$ and $n_{\mathrm{ext}}$, or $\epsilon_{\mathrm{e}}$ and $\epsilon_{\mathrm{B}}$. Some other parameters are anti-correlated, like $E_{\mathrm{0,iso}}^{\mathrm{c}}$ and $\epsilon_{\mathrm{B}}$, or $n_{\mathrm{ext}}$ and $\epsilon_{\mathrm{B}}$. Such degeneracies in the model parameters are expected when only the synchrotron component is observed (see e.g. Aksulu et al., 2022). In the case of GW 170817, the afterglow light curves can in first approximation be described by five quantities – its peak time, peak flux, rising and decreasing phase temporal slopes and spectral slope – which does not allow to constrain all the free parameters listed in Table 3.
As a post-processing step, we re-sample 20 0002000020\,00020 000 light curves at several observing frequencies ($3$ GHz, $1$ keV and $1$ TeV) and spectra at several observing times ($20$ days, $110$ days, $400$ days), where the set of parameters is directly drawn from the posterior sample, for each of the three fits. At each observing time (for the light curves) or at each frequency (for the spectra), we then determine the $68\%$ and $97.5\%$ credible intervals for the distribution of predicted flux densities, that we use to draw confidence contours around the median value of the light curves and spectra. Because the posterior samples are similar between the three models, we show only the results for the most realistic case (”SSC (with KN)” model) in Figs. 7 and 8 (radio and VHE light curve), and in Fig. 9 (spectrum at 110 days, close to the peak).
Figure 6: Posterior joint and marginalized distributions of the model’s free parameters, $E_{\mathrm{0,iso}}^{\mathrm{c}}$, $\theta_{\mathrm{c}}$, $\theta_{\mathrm{v}}$, $n_{\mathrm{ext}}$, $\epsilon_{\mathrm{B}}$, $\epsilon_{\mathrm{e}}$, $\zeta$, $p$ $a$, $\Gamma_{0}^{\mathrm{c}}$, $b$, for the ”SSC (with KN)” fit of the afterglow of GW 170817 presented in Section 4 which includes SSC diffusions in both Thomson and KN regimes. The priors are all uniform or log-uniform and are shown in Table 3. Fitted data includes all points until $t_{\rm max}=400\leavevmode\nobreak\ \mathrm{days}$, as well as $5$ early-time upper limits (see Section 4.1). The colored contours correspond to the $1\sigma$, $2\sigma$ and $3\sigma$ confidence intervals for each parameter.
4.3 Discussion: inferred parameters
As seen in Fig. 7, the afterglow of GW 170817 is very well fitted by the model. The model predicts accurately the flux at all times, with a very small dispersion. The dispersion of the predicted flux is of course larger at very early or very late times where we do not include any data point in the fit. The same quality of the model prediction is observed in the spectrum on Fig. 9. The frequency $\nu_{\mathrm{m,obs}}$ is below the data points at the lowest frequency, with some dispersion in absence of observational constraints. The critical frequency $\nu_{\mathrm{c,obs}}$ is found to be in X-rays, just above the highest frequency data points.
The inferred values of most model parameters (Fig. 5 and Table 5) are very similar to the results obtained in previous studies modelling the same afterglow with the synchrotron radiation from a decelerating structured jet (Troja et al., 2019; Lamb et al., 2019; Ghirlanda et al., 2019; Ryan et al., 2020), with some exceptions, due to different priors. A first difference concerns the viewing and core jet opening angles $\theta_{\mathrm{v}}$ and $\theta_{\mathrm{c}}$. Even if the values inferred by the ”SSC (with KN)” fit are comparable with values found in previous studies, the corresponding ratio $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}\sim 11$ (Table 6) is on the higher end of the other predictions. Nakar & Piran (2021) give a compilation of these predictions and show that afterglow light curves can only constrain this ratio and not the two parameters independently, which is confirmed by the correlation observed in Fig. 6. This difference is probably due to the flat and broad priors we use for these two parameters, whereas most previous studies use much stricter priors based on an external constraint, either on the viewing angle derived from GW data using the value of the Hubble constant from Planck (as in Troja et al., 2019; Ryan et al., 2020) or a constraint on the viewing angle and the core jet using VLBI imagery (as in Ghirlanda et al., 2019; Lamb et al., 2019). The marginalized distributions of $\theta_{\mathrm{v}}$ and $\theta_{\mathrm{c}}$ show that a restricted region of the parameter space has been explored by the MCMC chains for these parameters, as is also visible on Fig. 16.
Figure 7: Posterior distribution of the radio light curves at $\nu_{\rm obs}=3\leavevmode\nobreak\ \mathrm{GHz}$ for the ”SSC (with KN)” fit of the afterglow of GW 170817. Solid lines represent the median value at each observing time, dark contours the $68\%$ confidence interval and light contours the $97.5\%$ confidence interval. Late-time observations not used in the fit are shown in blue.
When comparing with the results of Ryan et al. (2020), where the assumptions are the closest to those of our ”no SSC” fit and where the light curves are computed using afterglowpy with which we compared our model in Section 3, we also find that our predicted values for $E_{0\mathrm{,iso}}^{\mathrm{c}}$ and $n_{\mathrm{ext}}$ are about one order of magnitude higher. From Fig. 6, it appears that it is also an effect of our different priors for the angles. However we observe the expected strong correlation between these two parameters and the inferred ratio $E^{\mathrm{c}}_{\mathrm{0,iso}}/n_{\mathrm{ext}}\sim 10^{55}\leavevmode\nobreak\ \mathrm{erg\cdot cm^{-3}}$ (Table 6) is close to the value obtained by Ryan et al. (2020). Using similar priors for the angles, or including a constraint on the external density from the observation of the host galaxy (Hallinan et al., 2017; Hajela et al., 2019) would then reduce the inferred values for these two parameters. We discuss below the fact that detections in the VHE range would help to break this degeneracy and allow for a more precise determination of the energy and external density independently of such external constraints.
@@ 4.3 Discussion: inferred parameters
The inferred values for the microphysics parameters $\epsilon_{\mathrm{B}}$ and $p$ are very close to those obtained in previous studies. The value of $\epsilon_{\mathrm{e}}$ is slightly lower, but we also find that the median value for the fraction of accelerated electrons is $\sim 0.15$, whereas this parameter is not included in past studies. Both parameters are strongly correlated, as seen in Fig. 6. Overall, the values of these microphysics parameters, including $\zeta$, are in a good agreement with the current understanding of the plasma physics at work in relativistic collisionless shocks (see e.g. Sironi et al., 2015) and comparable to values obtained for cosmological short GRBs by Fong et al. (2015). The inferred value $\zeta<1$ leads to the question of a possible contribution to the radiation of the remaining thermal electrons (see e.g. Warren et al., 2022). Assuming $\zeta=1$ as in previous studies is relevant in the case of GW 170817 as the observed emission is dominated by the synchrotron radiation, and the cooling break $\nu_{\mathrm{c,obs}}$ is not detected. However these microphysics parameter impact the synchrotron and SSC components in different manners, as discussed below, and must therefore be considered to predict the VHE emission.
Figure 8: Posterior distribution of the VHE light curves at $h\nu_{\rm obs}=1\leavevmode\nobreak\ \mathrm{TeV}$ for the ”SSC (with KN)” fit of the afterglow of GW 170817. Same conventions as in Fig. 7. Note that this is a prediction of the model, with no observational constraint included in the fit.
@@
Figure 9: Posterior distribution of the afterglow spectrum around its peak ($t_{\mathrm{obs}}=110\leavevmode\nobreak\ \mathrm{days}$) for the ”SSC (with KN)” fit of the afterglow of GW 170817. Data points show the multi-wavelength observations at $t_{\mathrm{obs}}\pm 4\leavevmode\nobreak\ \mathrm{days}$. The upper limit from H.E.S.S. is also indicated (Abdalla et al., 2020). The low-energy component (solid line) is produced by synchrotron radiation, while the high-energy emission (dashed line) is powered by SSC diffusions. Thick lines represent the median value at each observing frequency, dark contours the $68\%$ confidence interval and light contours the $97.5\%$ confidence interval. Some instrument observing spectral ranges are shown in colors.
@@
As shown in the second row of Fig. 5 and in Table 6, the inferred values of the parameters describing the initial lateral structure of the ejecta are in good agreement with the conditions established in Beniamini et al. (2020b) and listed in § 2.2 to get a single-peak light curve: note that in practice, double-peak light curves are not excluded a priori in our MCMC, but are disfavored due to the inclusion of early upper limits (Table 4). Finally, we note that the initial distribution of the Lorentz factor ($\Gamma^{\mathrm{c}}_{0}$ and $b$) is not well constrained, which was expected as most of the observed emission is produced in the self-similar stage of the deceleration. A correlation is observed between $\Gamma_{0}^{\mathrm{c}}$ and $b$. This can be explained by the fact that the lateral structure is required to reproduce the early-time slow rise of the light curve (Nakar & Piran, 2018). Specifically, some material needs to have lower initial Lorentz factors to overcome effects of Doppler boosting out of the line of sight. With a higher core $\Gamma_{0}^{\mathrm{c}}$, a steeper decrease in $\Gamma_{0}(\theta)$ is required for material at a given latitude $\theta$ to have a sufficiently low initial Lorentz factor that the observed slow rise can be reproduced, leading to higher values of $b$.

### v2

@@ 4.1 Observational data
In its numerical implementation, our model allows for simultaneous multi-wavelength modelling. We used $94$ data points, from radio to X-rays between $9.2$ and $380$ days [^3] Data points are provided by K. Mooley on github.com/kmooley/GW170817/, as compiled and reprocessed homogeneously by Makhathini et al. (2021). We also included the five constraining early-time upper limits also reported in Makhathini et al. (2021), as shown in Table 4. Early-time constraints are useful to restrict the posterior models to those with a single peak in flux density, as discussed in Sect. 2.2. All these flux measurements and upper limits were fitted simultaneously.
The likelihood that we used for our study is based on a modified $\chi^{2}$ calculation to account for the upper limits,
@@
where the first sum concerns observational data (subscripts $i$), while the second sum concerns upper limits (subscripts $j$). The quantities $m_{i}$ and $m_{j}$ are the model flux densities at each observing time and frequency of the observational data set, $d_{i}$ is the flux density value of a detection, $\sigma_{i}$ is its associated uncertainty, $u_{j}$ is the upper flux limit presented in Table 4, and $\sigma_{j}$ is an equivalent uncertainty that we arbitrarily defined as $\sigma_{j}=0.2u_{j}$, which roughly corresponds to the typical uncertainties for the detections. Adding this element accounts for some models whose predictions are slightly above the reported upper limit. The quantity $\max\{u_{j};m_{j}\}$ in the second term implies that it behaves as a penalty that is only added for predicted fluxes $m_{j}$ above the observed upper limits. Another option would be to force the likelihood to diverge as soon as one of the modelled fluxes is above the corresponding upper limit.
We chose to exclude late-time observations after $400\leavevmode\nobreak\ \mathrm{days}$ in the fits. We expect the lateral spreading of the jet to play a role at late times, when the shock front reaches Lorentz factors $\Gamma\lesssim 3$, by which point our model no longer accurately describes the jet dynamics. In addition, other emission sites (e.g. the kilonova afterglow; Nakar & Piran 2011) could also contribute to the observed flux at late times. It is also difficult to include the latest observations (Hajela et al. 2022), which were not included in the compilation by Makhathini et al. (2021): Late-time data in X-rays consist of only a few photons, and a correct modelling in this case should be expressed in photon counts and should account for the instrumental response and low-statistics effects. The result of such an analysis is currently unclear, with an uncertain late rebrightening (Hajela et al. 2022) (see the discussion in Troja et al. 2022). An interpretation of the late evolution of the afterglow of GW 170817 appears to require a dedicated study, and we therefore excluded this phase from the analysis presented here. Finally, we did not include the H.E.S.S. upper limit at the peak (Abdalla et al. 2020), or the Fermi large area telescope (LAT) and H.E.S.S. early upper limits (Ajello et al. 2018; Abdalla et al. 2017), as they do not constrain the fit (see Sect. 4.4).
@@
 [^4] The times are given in days, the frequencies in Hz, and the upper limits in $\mu\mathrm{Jy}$. These upper flux limits are selected as the most constraining of all the early observations listed in the compilation by Makhathini et al. (2021).
Figure 5: Comparison of the marginalised posterior distributions for the three fits of the afterglow of GW 170817 presented in Sect. 4, which differ only by the treatment of the radiative processes: ”no SSC” in blue; ”SSC (Thomson)” in green; ”SSC (with KN)” in orange. The first two rows show the inferred distributions of the free model parameters. The last row shows the distributions of several quantities derived from these parameters: The true energy of the jet $E_{0}$ (see Eq. (9)), the ratio $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}$, and the four quantities used in the conditions for single-peaked light curves from Beniamini et al. (2020b) and described in Sect. 2.2, $\log{\left(\theta_{\mathrm{c}}\Gamma_{0}^{\mathrm{c}}\right)}$, $8b/a$, $-b\frac{\log\theta_{\mathrm{c}}}{\log\Gamma_{0}^{\mathrm{c}}}$, and $(b-1)\frac{\log{\left(\theta_{\mathrm{v}}/\theta_{\mathrm{c}}\right)}}{\log{\left(\Gamma_{0}^{\mathrm{c}}\theta_{\mathrm{c}}\right)}}$.
@@ 4.2 Results from the afterglow fitting
Table 3: Free parameters of the MCMC and their prior bounds and shape.
@@
We performed a Bayesian analysis of the afterglow data of GW 170817 using the Markov chain Monte Carlo (MCMC) algorithm of the Python suite emcee (Foreman-Mackey et al. 2013, 2019). We ran three separate fits that differed only in the assumptions for the radiative processes in the shocked region.
  1. 1.
    ”no SSC”: Emission is only produced by synchrotron radiation, as described in Sari et al. (1998) and Sect. 2.4.4.
  2. 2.
    ”SSC (Thomson)”: SSC is taken into account assuming that all scatterings occur in the Thomson regime, as described in Sari & Esin (2001) and Sect. 2.4.5.
  3. 3.
    ”SSC (with KN)”: SSC is taken into account including the KN regime, as described in Nakar et al. (2009) and Sect. 2.4.6.
The priors used for these three fits are identical and presented in Table 3. We used (log-)uniform priors over large intervals in order to remain as agnostic as possible. Their bounds are shown in Table 3. Additional constraints on the viewing angle can be obtained by combining the GW signal and the accurate distance and localisation of the host galaxy (Finstad et al. 2018), or by combining the afterglow photometry and VLBI imagery (Govreen-Segal & Nakar 2023), and the density of the external medium can be constrained from direct observation of the host galaxy (Hallinan et al. 2017; Hajela et al. 2019). We did not include them to focus on the constraints obtained from the afterglow modelling alone.
For each of the three fits, we initialised $50$ chains and ran 10 0001000010\,00010 000 iterations per chain. After studying the convergence speed, we removed the first 2 00020002\,0002 000 iterations for each chain. We show the posterior distributions for the remaining 400 000400000400\,000400 000 parameter samples.
We present the marginalised posterior distribution of the free parameters for the three fits in Fig. 5, complemented by Fig. 16, which shows the corresponding joint and marginalised posterior distributions at $3\sigma$ credibility (corner plot). We also show in Fig. 5 the distribution of some derived quantities, the true initial kinetic energy of the ejecta $E_{0}$ as deduced from Eq. (9), the ratio of the viewing angle over the opening angle of the core jet $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}$, and the four conditions for single-peaked light curves described by Beniamini et al. (2020b) and in Sect. 2.2. The three fits lead to similar distributions across the parameter space, with a few exceptions that are discussed in Sect. 4.3. The median parameter values as well as their 90% credible intervals are reported in Table 5. We finally show in Table 6 the median values of the derived quantities shown in Fig. 5.
Figure 6: Posterior joint and marginalised distributions of the free model parameters, $E_{\mathrm{0,iso}}^{\mathrm{c}}$, $\theta_{\mathrm{c}}$, $\theta_{\mathrm{v}}$, $n_{\mathrm{ext}}$, $\epsilon_{\mathrm{B}}$, $\epsilon_{\mathrm{e}}$, $\zeta$, $p$ $a$, $\Gamma_{0}^{\mathrm{c}}$, $b$, for the ”SSC (with KN)” fit of the afterglow of GW 170817 presented in Sect. 4, which includes SSC scatterings in the Thomson and KN regimes. The priors are all uniform or log-uniform and are shown in Table 3. The fitted data include all points until $t_{\rm max}=400\leavevmode\nobreak\ \mathrm{days}$, as well as five early-time upper limits (see Sect. 4.1). The coloured contours correspond to the $1\sigma$, $2\sigma$, and $3\sigma$ confidence intervals for each parameter.
Table 4: Posterior values for the free parameters of the three fits of the afterglow of GW 170817.
@@
555 The three fits are presented in Sect. 4.2. The table reports the median value of the posterior distribution of each parameter. The reported uncertainties are the 90% credible intervals.
Figure 7: Posterior distribution of the radio light curves at $\nu_{\rm obs}=3\leavevmode\nobreak\ \mathrm{GHz}$ for the ”SSC (with KN)” fit of the afterglow of GW 170817. The solid lines represent the median value at each observing time, the dark contours are the $68\%$ confidence interval, and the light contours show the $97.5\%$ confidence interval. Late-time observations that were not used in the fit are shown in blue.
Figure 8: Posterior distribution of the VHE light curves at $h\nu_{\rm obs}=1\leavevmode\nobreak\ \mathrm{TeV}$ for the ”SSC (with KN)” fit of the afterglow of GW 170817. Same as in Fig. 7. This is a prediction of the model. No observational constraint is included in the fit.
The detailed joint posterior distributions for the ”SSC (with KN)” model where emission is produced by synchrotron radiation and SSC scatterings in the Thomson and KN regimes are shown in Fig. 6. In this figure, which is representative of the results obtained with the three fits, we observe typical correlations between some of the afterglow parameters, such as $\theta_{\mathrm{c}}$ and $\theta_{\mathrm{v}}$, $E_{\mathrm{0,iso}}^{\mathrm{c}}$ and $n_{\mathrm{ext}}$, or $\epsilon_{\mathrm{e}}$ and $\epsilon_{\mathrm{B}}$. Some other parameters are anti-correlated, such as $E_{\mathrm{0,iso}}^{\mathrm{c}}$ and $\epsilon_{\mathrm{B}}$, or $n_{\mathrm{ext}}$ and $\epsilon_{\mathrm{B}}$. These degeneracies in the model parameters are expected when only the synchrotron component is observed (see e.g. Aksulu et al. 2022). In the case of GW 170817, the afterglow light curves can to first approximation be described by five quantities: its peak time, peak flux, the temporal slopes of the rising and decreasing phase, and the spectral slope. This does not allow us to constrain all the free parameters listed in Table 3.
As a post-processing step, we resampled 20 0002000020\,00020 000 light curves at several observing frequencies ($3$ GHz, $1$ keV, and $1$ TeV) and spectra at several observing times ($20$ days, $110$ days, and $400$ days), where the set of parameters was directly drawn from the posterior sample for each of the three fits. At each observing time (for the light curves) or at each frequency (for the spectra), we then determined the $68\%$ and $97.5\%$ credible intervals for the distribution of predicted flux densities, which we used to draw confidence contours around the median value of the light curves and spectra. Because the posterior samples are similar for the three models, we only show the results for the most realistic case (”SSC (with KN)” model) in Figs. 7 and 8 (radio and VHE light curve), and in Fig. 9 (spectrum at 110 days, close to the peak).
4.3 Discussion: Inferred parameters
As shown in Fig. 7, the afterglow of GW 170817 is very well fitted by the model. The model accurately predicts the flux at all times, with a very small dispersion. The dispersion of the predicted flux is larger at very early or very late times, when no data point is included in the fit. The quality of the model prediction is the same for the spectrum in Fig. 9. The frequency $\nu_{\mathrm{m,obs}}$ is below the data points at the lowest frequency, with some dispersion in absence of observational constraints. The critical frequency $\nu_{\mathrm{c,obs}}$ is found to be in X-rays, just above the highest-frequency data points.
The inferred values of most model parameters (Fig. 5 and Table 5) are very similar to the results obtained in previous studies that modelled the same afterglow with the synchrotron radiation from a decelerating structured jet (Troja et al. 2019; Lamb et al. 2019; Ghirlanda et al. 2019; Ryan et al. 2020), with some exceptions that are due to different priors. A first difference concerns the viewing and core jet opening angles $\theta_{\mathrm{v}}$ and $\theta_{\mathrm{c}}$. Even though the values inferred by the ”SSC (with KN)” fit are comparable with values found in previous studies, the corresponding ratio $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}\sim 11$ (Table 6) is at the higher end of the other predictions. Nakar & Piran (2021) reported a compilation of these predictions and showed that afterglow light curves can only constrain this ratio and not the two parameters independently, which is confirmed by the correlation observed in Fig. 6. This difference is probably due to the flat and broad priors we used for these two parameters, whereas most previous studies used much stricter priors based on an external constraint, either on the viewing angle derived from GW data using the value of the Hubble constant from Planck (as in Troja et al. 2019; Ryan et al. 2020), or a constraint on the viewing angle and the core jet using VLBI imagery (as in Ghirlanda et al. 2019; Lamb et al. 2019). The marginalised distributions of $\theta_{\mathrm{v}}$ and $\theta_{\mathrm{c}}$ show that a restricted region of the parameter space has been explored by the MCMC chains for these parameters, as is also visible in Fig. 16.
When compared with the results of Ryan et al. (2020), whose assumptions are closest to those of our ”no SSC” fit and who computed the light curves using afterglowpy, with which we compared our model in Sect. 3, we also find that our predicted values for $E_{0\mathrm{,iso}}^{\mathrm{c}}$ and $n_{\mathrm{ext}}$ are higher by about one order of magnitude. Fig. 6 shows that it is also an effect of our different priors for the angles. However, we observe that the expected strong correlation between these two parameters and the inferred ratio $E^{\mathrm{c}}_{\mathrm{0,iso}}/n_{\mathrm{ext}}\sim 10^{55}\leavevmode\nobreak\ \mathrm{erg\cdot cm^{-3}}$ (Table 6) is close to the value obtained by Ryan et al. (2020). Using similar priors for the angles, or including a constraint on the external density from the observation of the host galaxy (Hallinan et al. 2017; Hajela et al. 2019), would then reduce the inferred values for these two parameters. We discuss below the fact that detections in the VHE range would help us to break this degeneracy and allow for a more precise determination of the energy and external density, independently of these external constraints.
@@ 4.3 Discussion: Inferred parameters
 [^6] The three fits are presented in Sect. 4.2 and the posterior values of the free parameters are reported in Table 5. The table reports the median value of posterior distribution of each quantity. The uncertainties reported are the 90% credible intervals. The four last rows correspond to the quantities that appear in the conditions for single-peak light curves (Sect. 2.2).
@@
The inferred values for the microphysics parameters $\epsilon_{\mathrm{B}}$ and $p$ are very close to those obtained in previous studies. The value of $\epsilon_{\mathrm{e}}$ is slightly lower, but we also find that the median value for the fraction of accelerated electrons is $\sim 0.15$, whereas this parameter is not included in past studies. Both parameters are strongly correlated, as shown in Fig. 6. Overall, the values of these microphysics parameters, including $\zeta$, agree well with the current understanding of the plasma physics at work in relativistic collisionless shocks (see e.g. Sironi et al. 2015), and they are comparable to the values obtained for cosmological short GRBs by Fong et al. (2015). The inferred value $\zeta<1$ leads to the question of a possible contribution to the radiation of the remaining thermal electrons (see e.g. Warren et al. 2022). Assuming $\zeta=1$ as in previous studies is relevant in the case of GW 170817 as the observed emission is dominated by the synchrotron radiation, and the cooling break $\nu_{\mathrm{c,obs}}$ is not detected. However, these microphysics parameters impact the synchrotron and SSC components in different manners, as discussed below, and they must therefore be considered to predict the VHE emission.
@@
Figure 9: Posterior distribution of the afterglow spectrum around its peak ($t_{\mathrm{obs}}=110\leavevmode\nobreak\ \mathrm{days}$) for the ”SSC (with KN)” fit of the afterglow of GW 170817. The data points show the multi-wavelength observations at $t_{\mathrm{obs}}\pm 4\leavevmode\nobreak\ \mathrm{days}$. The upper limit from H.E.S.S. is also indicated (Abdalla et al. 2020). The low-energy component (solid line) is produced by synchrotron radiation, while the high-energy emission (dashed line) is powered by SSC scattering. The thick lines represent the median value at each observing frequency, the dark contours show the $68\%$ confidence interval, and the light contours present the $97.5\%$ confidence interval. Some instrument observing spectral ranges are shown in colours.
@@
As shown in the second row of Fig. 5 and in Table 6, the inferred values of the parameters describing the initial lateral structure of the ejecta agree well with the conditions established in Beniamini et al. (2020b) and listed in Sect. 2.2 to obtain a single-peak light curve: In practice, double-peak light curves are not excluded a priori in our MCMC, but are disfavoured due to the inclusion of early upper limits (Table 4). Finally, we note that the initial distribution of the Lorentz factor ($\Gamma^{\mathrm{c}}_{0}$ and $b$) is poorly constrained, which was expected as most of the observed emission is produced in the self-similar stage of the deceleration. A correlation is observed between $\Gamma_{0}^{\mathrm{c}}$ and $b$. This can be explained by the fact that the lateral structure is required to reproduce the early-time slow rise of the light curve (Nakar & Piran 2018). Specifically, some material needs to have lower initial Lorentz factors to overcome effects of Doppler boosting out of the line of sight. With a higher core $\Gamma_{0}^{\mathrm{c}}$, a steeper decrease in $\Gamma_{0}(\theta)$ is required for material at a given latitude $\theta$ to have a sufficiently low initial Lorentz factor that the observed slow rise can be reproduced, leading to higher values of $b$.